# Large-scale distributed synchronization systems, using a cancel-on-completion redundancy mechanism

### v1

@@ Abstract
We consider a class of multi-agent distributed synchronization systems, which are modeled as $n$ particles moving on the real line. This class generalizes the model of a multi-server queueing system, considered in [15], employing so-called cancel-on-completion (c.o.c.) redundancy mechanism, but is motivated by other applications as well. In the multi-server queueing system a particle location represents a server workload. Under c.o.c. mechanism, when a job of class $j$ arrives, it selects $d_{j}$ particles uniformly at random, which try to jump forward, by random distances, but their advance is truncated at the new location of the $k_{j}$-th left-most particle ($k_{j}\leq d_{j}$). Between jumps all particles move to the left at constant speed, but cannot cross point 00 (workload cannot be less than 00). Thus, the multi-server queueing system is modeled as a particle system, regulated at the left boundary point.
The more general model of this paper is such that particles evolve the same way as the in left-regulated system, but we allow regulation boundaries on either side, or both sides, or no regulation at all. We consider the mean-field asymptotic regime, when the number of particles $n$ and the job arrival rates go to infinity, while the job arrival rates per particle remain constant. The results include: the existence/uniqueness of fixed points of mean-field limits (ML), which describe the limiting dynamics of the system; conditions for the steady-state asymptotic independence (concentration, as $n\to\infty$, of the stationary distribution on a single state, which is necessarily an ML fixed point); the limits, as $n\to\infty$, of the average velocity at which unregulated (free) particle system advances. In particular, our results for the left-regulated system unify and generalize the corresponding results in [15]. Our technical development is such that the systems with different types of regulation are analyzed within a unified framework. In particular, these systems are used as tools for analysis of each other.
@@ 1 Introduction
We consider a class of particle systems which generalizes the class of multi-server queueing models, considered in [15], with so-called cancel-on-completion (c.o.c.) redundancy. The evolution of server workloads in a multi-server system with c.o.c. redundancy can be viewed as evolution of particles on the real line, with regulation at the left boundary at 00 – particles (server workloads) are not allowed to move to the left of 00. The more general model of this paper is such that, in addition to left-regulated particle systems, we also consider right- and/or left-regulated, as well as unregulated (free) systems. We study the mean-field asymptotic regime, when the number of particles and the job arrival rates go to infinity, while the job arrival rates per particle remain constant. We obtain results on both the asymptotics of stationary distributions and on the behavior of mean-field limits, which describe the limiting dynamics of the system. (Unregulated systems were considered already in [15], but not their asymptotic behavior.) A feature of our technical development is that the systems with different types of regulation are analyzed within a unified framework. In particular, such systems are used as tools for analysis of each other.
@@
One can view the process described above as a particle system, if each server is associated with a “particle,” and its workload is interpreted as the particle “location” on the real line. With this view, an arriving class-$j$ “job” selects $d_{j}$ particles uniformly at random, and their random non-negative potential jump sizes are drawn from distribution $F_{j}$, which determine their potential new locations after the jumps. However, the actual new locations the selected particles jump to, are “truncated” by the $k_{j}$-th smallest potential new location (due to the c.o.c. mechanism). Between “job” arrivals, each particle moves to the “left” at constant speed $v=1$, but cannot move to left of the “regulation” boundary at 00; a particle system with this feature will be referred to as left-regulated. A system state is the empirical distribution of particles’ locations. Due to the jump truncation feature (c.o.c. mechanism) the system is “non-work-conserving” in the sense that the expected total displacement of the selected particles after a class-$j$ job arrival is not constant – it, in general, depends on the system state at the time of arrival, namely on the relative locations of selected particles. The non-work-conservation is the model feature which substantially complicates the system analysis (see section 1.1 in [15]).
@@
We study the mean-field asymptotic regime, when $n\to\infty$, $\lambda^{n}\to\lambda$, and $\pi^{n}_{j}\to\pi_{j}$ for all $j$. In this regime, deterministic trajectories arising as limits of the process dynamics are called mean-field limits (ML). We say that the steady-state asymptotic independence (SSAI) holds when the sequence of stationary distributions converges to the distribution concentrated on a single state (which is then necessarily an ML fixed point). Some of the main questions in this asymptotic regime concern the existence and characterization of ML fixed points and conditions under which the SSAI holds. Paper [15] focuses on these questions for the left-regulated system.
@@
In this paper we study particle systems with the same job arrival and jump truncation (c.o.c.) mechanism, under the same asymptotic regime, but more general in that we allow systems regulated on the left and/or right, as well as unregulated (free) systems. It is also convenient to have the speed $v$, at which each particle moves to the left, as an additional parameter, and consider the dependence of the system properties on $v$, for fixed parameters $\lambda$, $\{\pi_{j}\}$, and $\{F_{j}\}$.
@@
Additional question of interest for a free system is: If $v_{n}$ is the average velocity at which the entire particle system (say, its mean) advances to the “right,” does $\lim_{n\to\infty}v_{n}$ exist? If so, is the limit determined by the speed parameter of the free system ML fixed points (if any). Note that in a free system there is the obvious relation between ML fixed points and MLs that are traveling waves. Namely, for any $u\geq 0$, an ML fixed point for speed parameter $v$ corresponds to a traveling wave ML moving right with velocity $u$, when the speed parameter is $v-u$.
@@
Some of our main results (stated informally) are as follows. They hold under the assumption of finite second moment of job component sizes and very mild non-degeneracy assumptions.
  • For the free system, a unique ML fixed point exists for any speed $v$ within some non-empty range $[v_{min},v_{max}]$. (Theorem 7.4(i))
@@
While results proving SSAI for redundancy models are scarce, the SSAI conjecture is often employed to obtain estimates of the system performance metrics when the number of servers is large (cf. [17, 9, 12]).
The unregulated (free) version of our model is the system of particles evolving on the real line, without any boundaries. There is an extensive literature on the mean-field asymptotics for such systems, including mean-field limits, their traveling wave solutions (ML fixed points in our terminology), and SSAI. Cf. reviews in [16, 5].
@@ 2 Basic notation and terminology
We denote by $\mathbb{R}$ and $\mathbb{R}_{+}$ the sets of real and real non-negative numbers, respectively, and by $\mathbb{R}^{n}$ and $\mathbb{R}_{+}^{n}$ the corresponding $n$-dimensional product sets. By $\bar{\mathbb{R}}\doteq\mathbb{R}\cup\{\infty\}\cup\{-\infty\}$ we denote the two-point compactification of $\mathbb{R}$, where $\infty$ and $\infty$- ∞ are the points at infinity and minus infinity, with the natural topology and a consistent with it metric, so that $\bar{\mathbb{R}}$ [and $\mathbb{R}$] is complete and separable. Analogously, $\bar{\mathbb{R}}_{+}\doteq\mathbb{R}_{+}\cup\{\infty\}$ is the one-point compactification of $\mathbb{R}_{+}$. For a vector $q=(q_{i})\in\mathbb{R}^{n}$, $|q|\doteq\sum_{i}|q_{i}|$. Inequalities applied to vectors [resp. functions] are understood component-wise [resp. for every value of the function argument]. We say that a function is RCLL [resp. LCRL] if it is right-continuous with left-limits [resp. left-continuous with right-limits]. A scalar function $f(w)$ is called $c$-Lipschitz if it is Lipschitz continuous with constant $c$. The sup-norm of a scalar function $f(w)$ is denoted $\|f(\cdot)\|\doteq\sup_{w}|f(w)|$; the corresponding convergence is denoted by $\stackrel{{\scriptstyle\|\cdot\|}}{{\longrightarrow}}$. U.o.c. convergence means uniform on compact sets convergence, and is denoted by $\stackrel{{\scriptstyle u.o.c.}}{{\longrightarrow}}$. We use notation: $a\vee b\doteq\max\{a,b\}$, $a\wedge b\doteq\min\{a,b\}$. Abbreviation w.r.t. means with respect to; a.e. means almost everywhere w.r.t. Lebesgue measure; WLOG means without loss of generality; RHS and LHS means right-hand side and left-hand side, respectively.
@@
We will equip the space $\tilde{\cal X}$ with the topology of weak convergence of the corresponding distributions on $\bar{\mathbb{R}}$; equivalently, $y\to x$ if and only if $y_{w}\to x_{w}$ for each $w\in(-\infty,\infty)$ where $x$ is continuous. Furthermore, on $\tilde{\cal X}$ we consider the following metric $L$, consistent with the weak convergence topology. We map an element $x\in\tilde{\cal X}$ into the (proper or non-proper) distribution function $\phi=\phi(x)$ on $\mathbb{R}$, where $\phi_{s}=1-x_{-\log(1-s)}$ for $10\leq s<$ 1, $\phi_{s}=1-x_{\log(1+s)}$ for $-1\leq s<0$, $\phi_{s}=0$ for $s<-1$, $\phi_{s}=1$ for $s\geq 1$. Then $L(x,y)\doteq\hat{L}(\phi(x),\phi(y))$, where $\hat{L}$ is the Levy-Prohorov metric (cf. [8]). It is easy to see that the convergence in metric $L$ is indeed equivalent to the weak convergence. Clearly, $\tilde{\cal X}$ is compact. Space ${\cal X}$ inherits the topology and metric of space $\tilde{\cal X}$ – for these purposes we view any $x\in{\cal X}$ as an element of $\tilde{\cal X}$ via the convention that $x_{w}=1$ for $w<0$. Clearly, ${\cal X}$ is also compact.
@@ 3.1 Multi-server queueing system with cancel-on-completion redundancy
A property of the c.o.c. redundancy which complicates its analysis is that it is non-work-conserving in the following sense. Suppose a class $j$ job arrives. Then, in general, the average total amount of new workload added by the arriving job to the selected servers depends not only on the distribution $F_{j}$ of the component sizes, but also on the current workloads of the selected servers or, more specifically, on their relative workloads (i.e. workload differences). In particular, the average rate per server at which new workload is added to the system is not known a priori. (If for each class $j$ either $k_{j}=d_{j}$ or the component sizes are i.i.d. exponential, the work-conservation does hold – the average total amount of new workload added by an arriving class $j$ job depends on $j$ only; see [15, 14].)
In this paper we will primarily use the following “particle” language to describe the system dynamics. Namely, we will identify each server with a “particle,” the server workload with “particle location,” and the workload evolution with “particle movement.” The basic dynamics of a particle (server workload) is as follows. If/when a job arrival adds workload $\kappa\geq 0$ to the server, the particle jumps “right” by $\kappa$. Between job arrivals, the particle (workload) moves “left” at the constant speed $1$ until and unless it “hits” the left “regulation boundary” at 00; if the particle does hit the boundary 00, it stays at it until and unless it jumps right due to a job arrival.
@@ 3.2 More general class of systems
  • Instead of a fixed left boundary at 00, the system may have: both left boundary at point $A$ and right boundary at point $B\geq A$; or only one – left or right – boundary; or have no boundaries at all. If there is a right boundary at point $B$, this means that if, after following the c.o.c. procedure upon a job arrival, some selected particles would “jump over” point $B$, they land exactly at $B$ instead.
System with finite left boundary $A$ will be called left-regulated, and without left boundary (which can be viewed as $A=-\infty$) – left-unregulated. The terms right-regulated, right-unregulated, both-sides-regulated, etc., will have analogous self-explanatory meanings. Unregulated system – with neither left nor right boundary – will also be called free (as in [15]). The interval between the left and right boundaries $A$ and $B$ is sometimes called the system frame (as in [15]). For example, frame $(-\infty,0]$ means left-unregulated system with right boundary at 00; finite frame $[A,B]$ means both-sides-regulated system; and so on.
@@
For both applications and analysis, for an anchored system it is sometimes convenient to consider its unanchored version, where particles do not move left between jumps, but instead the boundaries (if any) move right at the constant speed $v$. Clearly, anchored and the corresponding unanchored system are equivalent, in the sense that the former is the latter viewed in the system of coordinates moving right at constant speed $v$. (In other words, the unanchored system has a frame moving right at speed $v$, and the anchored version is obtained by considering the system evolution with respect to its frame.)
@@ 3.3 Assumptions on job component size distributions
$\mathbb{E}\xi^{(j)}_{1}<\infty~{}~{}\mbox{(and then $\mathbb{E}\sum_{i=1}^{d_{j}}\xi^{(j)}_{i}=d_{j}\mathbb{E}\xi^{(j)}_{1}<\infty$),}$, (1)
@@
That is, $x^{n}_{w}(t)$ is the fraction of particles $i$ with $W_{i}^{n}(t)>w$, i.e. those located to the right of point $w$. Then $x^{n}(t)=(x^{n}_{w}(t),~{}w\geq 0)$ is a projection of $W^{n}(t)$, and $x^{n}(t),~{}t\geq 0,$ is also a Markov process.
@@
In the special case of a regulated system, the process has simple structure – it is irreducible, possessing a fixed state being a renewal atom. For example, suppose the system is left-regulated at boundary point 00. (It also may or may not have a right boundary.) Then the state $(0,\ldots,0)$ of $W^{n}(\cdot)$, with all particles at the left boundary, it a renewal atom; this state we will call “empty”, because in terms of the multi-server queueing system corresponds to all servers being idle. The corresponding renewal atom for the process $x^{n}(\cdot)$ is the state $x^{\emptyset}$, defined by
@@
it will also be call the empty state. The irreducibility of $W^{n}(\cdot)$ [resp., $x^{n}(\cdot)$] follows from the facts that the empty state is reachable from any other, and that the job arrival process is Poisson.
@@
It corresponds to all particles located at +$\infty$+ ∞.

### v2

@@ Abstract
We consider a class of multi-agent distributed synchronization systems, which are modeled as $n$ particles moving on the real line. This class generalizes the model of a multi-server queueing system, considered in [15], employing so-called cancel-on-completion (c.o.c.) redundancy mechanism, but is motivated by other applications as well. In the multi-server queueing system a particle location represents a server workload. Under c.o.c. mechanism, when a job of class $j$ arrives, it selects $d_{j}$ particles uniformly at random, which try to jump forward, by random distances, but their advance is truncated at the new location of the $k_{j}$-th left-most selected particle ($k_{j}\leq d_{j}$). Between jumps all particles move to the left at constant speed, but cannot cross point 0 (workload cannot be less than 0). Thus, the multi-server queueing system is modeled as a particle system, regulated at the left boundary point.
The more general model of this paper is such that particles evolve the same way as the in left-regulated system, but we allow regulation boundaries on either side, or both sides, or no regulation at all. We consider the mean-field asymptotic regime, when the number of particles $n$ and the job arrival rates go to infinity, while the job arrival rates per particle remain constant. The system state for a given $n$ is the empirical distribution of the particles’ locations. Our results include: the existence/uniqueness of fixed points of mean-field limits (ML), which describe the limiting dynamics of the system; conditions for the steady-state asymptotic independence (concentration, as $n\to\infty$, of the stationary distribution on a single state, which is necessarily an ML fixed point); the limits, as $n\to\infty$, of the average velocity at which unregulated (free) particle system advances. In particular, our results for the left-regulated system unify and generalize the corresponding results in [15]. Our technical development is such that the systems with different types of regulation are analyzed within a unified framework. In particular, these systems are used as tools for analysis of each other.
@@ 1 Introduction
1.1 Model introduction and motivation
We consider a class of particle systems which generalizes the class of multi-server queueing models, considered in [15], with so-called cancel-on-completion (c.o.c.) redundancy. The evolution of server workloads in a multi-server system with c.o.c. redundancy can be viewed as evolution of particles on the real line, with regulation at the left boundary at 0 – particles (server workloads) are not allowed to move to the left of 0. The more general model of this paper is such that, in addition to left-regulated particle systems, we also consider right- and/or left-regulated, as well as unregulated (free) systems. We study the mean-field asymptotic regime, when the number of particles and the job arrival rates go to infinity, while the job arrival rates per particle remain constant. We obtain results on both the asymptotics of stationary distributions and on the behavior of mean-field limits, which describe the limiting dynamics of the system. (Unregulated systems were considered already in [15], but not their asymptotic behavior.) A feature of our technical development is that the systems with different types of regulation are analyzed within a unified framework. In particular, such systems are used as tools for analysis of each other.
@@ 1.1 Model introduction and motivation
One can view the process described above as a particle system, if each server is associated with a “particle,” and its workload is interpreted as the particle “location” on the real line. With this view, an arriving class-$j$ “job” selects $d_{j}$ particles uniformly at random, and their random non-negative potential jump sizes are drawn from distribution $F_{j}$, which determine their potential new locations after the jumps. However, the actual new locations the selected particles jump to, are “truncated” by the $k_{j}$-th smallest potential new location (due to the c.o.c. mechanism). Between “job” arrivals, each particle moves to the “left” at constant speed $v=1$, but cannot move to left of the “regulation” boundary at 0; a particle system with this feature will be referred to as left-regulated. Due to the jump truncation feature (c.o.c. mechanism) the system is “non-work-conserving” in the sense that the expected total displacement of the selected particles after a class-$j$ job arrival is not constant – it, in general, depends on the system state at the time of arrival, namely on the relative locations of selected particles. The non-work-conservation is the model feature which substantially complicates the system analysis (see section 1.1 in [15]).
@@
We study the mean-field asymptotic regime, when $n\to\infty$, $\lambda^{n}\to\lambda$, and $\pi^{n}_{j}\to\pi_{j}$ for all $j$. The system state for a given $n$ is the empirical distribution of particles’ locations. In this regime, deterministic trajectories arising as limits of the process dynamics are called mean-field limits (ML). We say that the steady-state asymptotic independence (SSAI) holds when the sequence of stationary distributions converges to the distribution concentrated on a single state (which is then necessarily an ML fixed point). Some of the main questions in this asymptotic regime concern the existence and characterization of ML fixed points and conditions under which the SSAI holds. Paper [15] focused on these questions for the left-regulated system.
@@
In this paper we study particle systems with the same job arrival and jump truncation (c.o.c.) mechanism, under the same asymptotic regime, but more general in that we allow systems regulated on the left and/or right, as well as unregulated (free) systems. It is also convenient to have the speed $v\geq 0$, at which each particle moves to the left, as an additional parameter, and consider the dependence of the system properties on $v$, for fixed parameters $\lambda$, $\{\pi_{j}\}$, and $\{F_{j}\}$.
@@
Additional question of interest for a free system is: If $v_{n}$ is the average velocity at which the entire particle system (say, its mean) advances to the “right,” does $\lim_{n\to\infty}v_{n}$ exist? If so, is the limit determined by the speed parameter of the free system ML fixed points (if any). Note that in a free system there is the obvious relation between ML fixed points and MLs that are traveling waves. Namely, for any $u\geq 0$, an ML fixed point for speed parameter $v$ corresponds to a traveling wave ML moving right with velocity $u$, when the speed parameter is vuv-u.
@@
1.2 Paper contributions
The main contributions of this paper are that we: (a) generalize and unify results for the left-regulated system, as compared to [15]; (b) obtain results for the free and left-regulated systems; and (c) our analysis demonstrates and exploits the connections between the systems with different types of regulation. Specifically, some of our main results, stated here informally, are as follows. (We give references to formal results, stated in Section 7.) The results hold under the assumption of finite second moment of job component sizes and very mild non-degeneracy assumptions.
  • For the free system, a unique ML fixed point exists for any speed $v$ within some non-empty range $v_{min},v_{max}]$. (Theorem 7.4(i))
@@ 1.3 Prior work
While results proving SSAI for redundancy models are scarce, the SSAI conjecture is often employed to compute the corresponding ML fixed point (on which the stationary distribution concentrates as $n\to\infty$, assuming SSAI holds), which is then used to obtain estimates of the system performance metrics when the number of servers is large (cf. [17, 9, 12]). Therefore, for systems, for which we prove SSAI in this paper, we rigorously substantiate such computations and estimates.
Recall that the unregulated (free) version of our model is the system of particles evolving on the real line, without any boundaries. There is an extensive literature on the mean-field asymptotics for such systems, including mean-field limits, their traveling wave solutions (ML fixed points in our terminology), and SSAI. Cf. reviews in [16, 5].
@@ 2 Basic notation and terminology
We denote by $\tilde{\cal X}$ [resp., $\cal X$}] the set of non-increasing RCLL functions $x=(x_{w},~w\in\mathbb{R})$, [resp., $x=(x_{w},~w\in\mathbb{R}_{+})$,] taking values in $[0,1]$. An element $x\in\tilde{\cal X}$ [resp., $x\in{\cal X}$] is naturally interpreted as complementary distribution function on $\bar{\mathbb{R}}$ [resp., $\bar{\mathbb{R}}_{+}$], with $1-x_{w}$ being the measure of $\infty,w] [$resp., $[0,w]$]. An element $x\in\tilde{\cal X}$ is proper if $x_{\infty}\doteq\lim_{w\to\infty}x_{w}=0$ and $x_{-\infty}\doteq\lim_{w\to-\infty}x_{w}=1$, and improper otherwise; in any case, $x_{\infty}$ and $1-x_{-\infty}$ give the measures of $\{+\infty\}$ and $\{-\infty\}$, respectively. Analogously, an element $x\in{\cal X}$ is proper if $x_{\infty}=0$ and and improper otherwise. (With some abuse of standard terminology, in this paper we sometimes refer to elements of $\tilde{\cal X}$ and $\cal X$} as distributions.) Denote by $\tilde{\cal X}^{pr}$ [resp., ${\cal X}^{pr}$] the subset of proper elements of $\tilde{\cal X}$ [resp., $\cal X$}].
We will equip the space $\tilde{\cal X}$ with the topology of weak convergence of the corresponding distributions on $\bar{\mathbb{R}}$; equivalently, $y\to x$ if and only if $y_{w}\to x_{w}$ for each $w\in(-\infty,\infty)$ where $x$ is continuous. Furthermore, on $\tilde{\cal X}$ we consider the following metric $L$, consistent with the weak convergence topology. We map an element $x\in\tilde{\cal X}$ into the (proper or non-proper) distribution function $\phi=\phi(x)$ on $\mathbb{R}$, where $\phi_{s}=1-x_{-\log(1-s)}$ for $0\leq s<1$, $\phi_{s}=1-x_{\log(1+s)}$ for $1\leq s<0$, $\phi_{s}=0$ for s<1s<-1, $\phi_{s}=1$ for $s\geq 1$. Then $L(x,y)\doteq\hat{L}(\phi(x),\phi(y))$, where $\hat{L}$ is the Levy-Prohorov metric (cf. [8]). It is easy to see that the convergence in metric $L$ is indeed equivalent to the weak convergence. Clearly, $\tilde{\cal X}$ is compact. Space $\cal X$} inherits the topology and metric of space $\tilde{\cal X}$ – for these purposes we view any $x\in{\cal X}$ as an element of $\tilde{\cal X}$ via the convention that $x_{w}=1$ for $w<0$. Clearly, $\cal X$} is also compact.
@@ 3.1 Multi-server queueing system with cancel-on-completion redundancy
Clearly, when a job arrives and selects a server, the actual amount $\kappa$ of workload added to the server may be smaller than the size $\xi$ of the job component placed on the server, i.e. $0\leq\kappa\leq\xi$, due to possible cancellation (of a part or all) of the component service. Moreover, $\kappa$ depends not only on the realization of the job component sizes, but also on the current workloads of the selected servers or, more specifically, on their relative workloads, i.e. workload differences. (Also note that due to FCFS discipline at each server, the actual amounts of workload added to selected servers do not depend on the future job arrivals; so that new workloads are well-defined after each new arrival.) Because of this, c.o.c. redundancy in general has a property, which substantially complicates its analysis, namely it is non-work-conserving in the following sense. Suppose a class $j$ job arrives. Then, in general, the average total amount of new workload added by the arriving job to the selected servers depends not only on the distribution $F_{j}$ of the component sizes, but also on the relative workloads of the selected servers. In particular, the average rate per server at which new workload is added to the system is not known a priori. (If for each class $j$ either $k_{j}=d_{j}$ or the component sizes are i.i.d. exponential, the work-conservation does hold – the average total amount of new workload added by an arriving class $j$ job depends on $j$ only; see [15, 14].)
In this paper we will primarily use the following “particle” language to describe the system dynamics. Namely, we will identify each server with a “particle,” the server workload with “particle location,” and the workload evolution with “particle movement.” The basic dynamics of a particle (server workload) is as follows. If/when a job arrival adds workload $\kappa\geq 0$ to the server, the particle jumps “right” by $\kappa$. Between job arrivals, the particle (workload) moves “left” at the constant speed $1$ until and unless it “hits” the left “regulation boundary” at 0; if the particle does hit the boundary 0, it stays at it until and unless it jumps right due to a job arrival.
@@ 3.2 More general class of systems
  • Instead of a fixed left boundary at 0, the system may have: both left boundary at point $A$ and right boundary at point $B\geq A$; or only one – left or right – boundary; or have no boundaries at all. If there is a right boundary at point $B$, this means that if, after following the c.o.c. procedure upon a job arrival, some selected particles would “jump over” point $B$, they land exactly at $B$ instead.
System with finite left boundary $A$ will be called left-regulated, and without left boundary (which can be viewed as $A=$-$\infty$) – left-unregulated. The terms right-regulated, right-unregulated, both-sides-regulated, etc., will have analogous self-explanatory meanings. Unregulated system – with neither left nor right boundary – will also be called free (as in [15]). The interval between the left and right boundaries $A$ and $B$ is sometimes called the system frame (as in [15]). For example, frame $(-\infty,0]$ means left-unregulated system with right boundary at 0; finite frame $[A,B]$ means both-sides-regulated system; and so on.
@@
For both applications and analysis, for an anchored system it is sometimes convenient to consider its unanchored version, where particles do not move left between jumps, but instead the boundaries (if any) move right at the constant speed $v$. Clearly, anchored and the corresponding unanchored system are equivalent, in the sense that the former is the latter viewed in the system of coordinates moving right at constant speed $v$. (In other words, the unanchored system has a frame moving right at speed $v$, and the anchored version is obtained by considering the system evolution with respect to its frame.) Obviously, in the special case of $v=0$ the anchored and unanchored systems are indistinguishable. The case of $v>0$ will of primary interest in this paper, but some of the results apply more generally to $v\geq 0$.
@@ 3.3 Assumptions on job component size distributions
$\mathbb{E}\xi^{(j)}_{1}<\infty~~\mbox{(and then $\mathbb{E}\sum_{i=1}^{d_{j}}\xi^{(j)}_{i}=d_{j}\mathbb{E}\xi^{(j)}_{1}<\infty$),}$ (1)
@@
That is, $x^{n}_{w}(t)$ is the fraction of particles $i$ with $W_{i}^{n}(t)>w$, i.e. those located to the right of point $w$. Then $x^{n}(t)=(x^{n}_{w}(t),~w\geq 0)$ describes the empirical distribution of particles’ locations, it is a projection of $W^{n}(t)$, and $x^{n}(t),~t\geq 0,$ is a Markov process.
@@
In the special case of a regulated system, the process has simple structure – it is irreducible, possessing a fixed state being a renewal atom. For example, suppose the system is left-regulated at boundary point 0. (It also may or may not have a right boundary.) Then the state $(0,\ldots,0)$ of $W^{n}(\cdot)$, with all particles at the left boundary, is a renewal atom; this state we will call “empty”, because in terms of the multi-server queueing system it corresponds to all servers being idle. The corresponding renewal atom for the process $x^{n}(\cdot)$ is the state $x^{\emptyset}$, defined by
@@
it will also be called the empty state. The irreducibility of $W^{n}(\cdot)$ [resp., $x^{n}(\cdot)$] follows from the facts that the empty state is reachable from any other, and that the job arrival process is Poisson.
@@
It corresponds to all particles located at $\infty$.